suppose that the short run production function of shoe manufactured by certain ethiopian firm is given byq=4lsquare-0.2lcube where q reprsents the anual quantity of shoe quantity of shoe produced find AP of labour and MPof labour





Find the value of L that maxmize the out put





Find the value of L that maxmize marginal product





Find the valurle of L that maxmize averge product

1
Expert's answer
2022-02-07T15:34:07-0500

AP of Labor and MP of Labor

"Q=4L^2-0.2L^3"

"AP(L)=\frac{Q}{L} =\frac{4L^2-0.2^3}{L}=4L-0.2L^2"

"MP(L)=\frac{\delta Q}{\delta L}= 8L-0.6L^2"


Value of L Maximizing Output

Here we take the value of MP and equate it to zero, and solve for L

"8L-0.6L^2=0"

"8L=0.6L^2"

"8=0.6L"

"\bold{L=13.33}"


Value of L that Maximize MP

Is taking the second derivative of the MP and equating it to zero.

"\frac{\delta MP}{\delta L}=8-1.2L"

"8-1.2L=0"

"\bold{L=6.67}"


Value of L that Maximizes Average Product

"MP=AP"

we find the derivative of AP and equate it to zero

"\frac{\delta AP}{\delta L} =4-0.4L=0"

"4=0.4L"

"\bold{L=10}"

At 10 units of labor, AP is maximized.